Multiply by 19: Multiply by 20 and subtract the original number.

Grade:11

1 Answers

Parvez ali
10 years ago
yes it is true

19*n = (20-1)*n
= 20*n-1*n
= 20*n-n



Thanks & Regards
Parvez Ali
askIITians Faculty
Btech
ISM Dhanbad